Dae 21 1 (1o markaj Distinguish between operating income a and net income 2 (20 marks) Ear etal Products reported bhe foilowing compones costs for Ape 29 in thousandsj Direct manufacturing labor salaries Manufacturing overhead Fringe benefits on direct manufacturing labor Production setup Other manufacturing overhead 60 240 Please note that Subcontracting Costs are treated as a direct cost item separate from drect materials Compute (plerime.costs and (bl the.conversion costs of EBY Metal Products using the cos classifications descrbed above. (422 t Direct.labe -loo
Answer #1

Answers :

1. Distinction between Operating Income and Net Income -

Operating Income Net Income
Also known as 'Earnings before interest and taxes'. Also known as the 'Profit'.
It is the income that a business generates from its operations. It is the final income that a business gets after considering all its expenses in a year.
Operating income is calculated by reducing operating expenses like wages and depreciation/ Ammortization from the Gross profit. Net income is calculated by adding back all other incomes like interest on investments etc and reducing all other expenses including taxes from Operating income.
Operating income provides the general view of business from its day to day operations. Net income provides the overall view of the entire business.
Operating income is not carried to the balance sheet or to next year as Retained earnings. Net Income is carried to the balance sheet and to the next year as Retained earnings.
A company's financial stability or position not depended on operating income. A company’s financial stability or position is depended on the net income.
Operating income is the earnings before tax and interests. Net income is the earnings after interest and taxes.

2.(a).Calculation of Prime cost-

Prime Cost =

Direct materials + Direct labor cost + Subcontracting cost

= 430 + 110 + 120 = $ 660

(b). Calculation of conversion cost

Conversion cost is the cost incurred to convert the raw materials into the form of finished goods .

= Direct Labor cost + Manufacturing overheads + subcontracting cost

= 110 +40 + 60 + 240 + 120 = $ 570